OpenStudy (anonymous):

express sqrt of 4x^2-9 in the form k sqrt x^2-a^2

jimthompson5910 (jim_thompson5910):

\[\large \sqrt{4x^2-9}\] \[\large \sqrt{4(x^2-\frac{9}{4})}\] \[\large \sqrt{4}\sqrt{x^2-\frac{9}{4}}\] \[\large 2\sqrt{x^2-\frac{9}{4}}\] \[\large 2\sqrt{x^2-(\frac{3}{2})^2}\] So \[\large \sqrt{4x^2-9}=2\sqrt{x^2-(\frac{3}{2})^2}\] Note: k=2 and a=3/2

OpenStudy (anonymous):

thank you